# Uncovering Treatment Effect Heterogeneity in Pragmatic Gerontology Trials

### v1

@@ Abstract
Detecting heterogeneity in treatment response enriches the interpretation of gerontologic trials. In aging research, estimating the effect of the intervention on clinically meaningful outcomes faces analytical challenges when it is truncated by death. For example, in the Whole Systems Demonstrator trial, a large cluster-randomized study evaluating telecare among older adults, the overall effect of the intervention on quality of life was found to be null. However, this marginal intervention estimate obscures potential heterogeneity of individuals responding to the intervention, particularly among those who survive to the end of follow-up. To explore this heterogeneity, we adopt a causal framework grounded in principal stratification, targeting the Survivor Average Causal Effect (SACE)—the treatment effect among “always-survivors,” or those who would survive regardless of treatment assignment. We extend this framework using Bayesian Additive Regression Trees (BART), a nonparametric machine learning method, to flexibly model both latent principal strata and stratum-specific potential outcomes. This enables the estimation of the Conditional SACE (CSACE), allowing us to uncover variation in treatment effects across subgroups defined by baseline characteristics. Our analysis reveals that despite the null average effect, some subgroups experience distinct quality of life benefits (or lack thereof) from telecare, highlighting opportunities for more personalized intervention strategies. This study demonstrates how embedding machine learning methods, such as BART, within a principled causal inference framework can offer deeper insights into trial data with complex features including truncation by death and clustering—key considerations in analyzing pragmatic gerontology trials.
Keywords: Whole Systems Demonstrator Telecare Questionnaire Study; gerontology trials, cluster randomized trial; pragmatic trials, principal stratification; machine learning; Bayesian inference; survivor average causal effect
@@ 1 Introduction
Pragmatic trials are a cornerstone of gerontologic research, providing rigorous yet policy-relevant evidence on interventions that aim to support independence, quality of life, and health in late life.roland1998understanding; resnick2022pragmatic Their strengths lie in large, heterogeneous populations and real-world delivery, often using cluster-randomized designs. eldridge2012practical However, these very characteristics can complicate the inference on the intervention performance: treatment effects averaged across diverse participants may appear modest or null, even when meaningful benefits (or harms) exist for specific subgroups. Conventional subgroup analyses, typically prespecified along a few baseline dimensions, are rarely sufficient to detect more complex forms of heterogeneity. As a result, valuable information may remain hidden, and seemingly “negative” trials risk being dismissed despite containing signals that could guide intervention targeting and future research. As a result, machine learning and other data-driven approaches are increasingly valuable for estimating individualized treatment effects and detecting complex, multidimensional patterns of heterogeneity that conventional subgroup analyses may overlook kent2018personalized; halvorsrud2007conceptualization. As one prominent example, parallel advances in machine learning for causal inference have shown that Bayesian Additive Regression Trees (BART) can flexibly learn complex regression surfaces with built-in regularization chipman2010bart, and have been recently applied to treatment effect heterogeneity detection in clinical trials. chen2024bayesian; hu2021estimating; ghazi2025treatment.
@@
Yet, in gerontologic trials, these opportunities come with unique complications. quality of life (QoL) outcomes that are among the most clinically meaningful endpoints in late-life intervention studies are a prime example. halvorsrud2007conceptualization; netuveli2008quality Measures such as EQ-5D or SF-12 are widely used to capture well-being, functional status, and independence—dimensions of aging that survival alone cannot reflect. group1990euroqol; ware199612 However, QoL outcomes face a special challenge: they are undefined for participants who die before follow-up, or who deteriorate so severely that reporting is impossible. This phenomenon, known as truncation by death, is especially common in gerontology trials, where mortality is non-negligible during extended follow-up. rubin2006causal Because participants who die may differ systematically from those who survive, and because interventions themselves may influence participant survival, analyses restricted to observed survivors risk introducing survivor bias and yield estimates that are difficult to interpret causally. One principled solution comes from principal stratification, which defines causal effects within strata of participants classified by their potential survival under both treatment and control.frangakis2002principal; zhang2003estimation In this framework, the survivor average causal effect (SACE) focuses on the “always-survivors”—those who would survive regardless of assignment—for whom both counterfactual QoL outcomes are well defined. Extending this idea to covariate conditional effects (CSACE) enables the investigation of how survivor responses vary across baseline characteristics, uncovering heterogeneity in a way that standard subgroup analyses cannot.chen2024bayesian
@@
The death truncation challenge is compounded by the fact that many pragmatic trials in gerontology adopt cluster randomization, either at the practice, facility, or community level, to reduce contamination, facilitate implementation, or reflect the way interventions are delivered in routine care. eldridge2012practical; murray1998design Such designs are especially common when evaluating service models, long-term care interventions, or technologies that naturally operate at a group level. For example, the Whole System Demonstrator trial of telecare randomized at the general practice level, steventon2012effect; hirani2014effect while cluster-randomized or cluster-stepped wedge designs have also been used in fall prevention, dementia care, and palliative interventions in older populations. (e.g., bhasin2020randomized; lamb2020screening Compared with individually randomized trials, CRTs introduce additional sources of correlation and heterogeneity, requiring methods that accounts for within-cluster dependence. A similar issue can arise in another type of closely related design in pragmatic trials called individually randomized group treatment trials (IRGTs), and a description of IRGT can be found elsewhere. tong2024designing Recent work has estimated SACE (and covariate conditional extensions, CSACE) in CRTs using parametric principal strata and outcome models with linear mixed effects and either Bayesian or EM-based estimation. tong2023bayesian; wang2024mixed While these approaches are interpretable and computationally convenient, they rely on linear specifications that may miss nonlinearities and high order interactions in rich baseline covariates, and they may be sensitive to functional-form misspecification.
@@
Another common complication in pragmatic gerontologic trials is missing data. Beyond death truncation, loss to follow-up is frequent as participants relocate, withdraw consent, or become unable to complete questionnaires due to cognitive or physical decline. little2012prevention; national2011prevention In such settings, outcomes may be missing even among survivors, creating additional uncertainty in treatment effect estimation. Moreover, survival status itself can sometimes be unobserved—for example, when participants move outside the health system or registry coverage—blurring the distinction between death-related truncation and standard missingness. chatfield2005systematic These challenges are especially pronounced in aging populations, where comorbidity, frailty, and mobility constraints increase the risk of incomplete follow-up chatfield2005systematic; agogo2018longitudinal. For the estimation of treatment effect heterogeneity, failure to account for missing data can distort subgroup contrasts or mask signals of effect modification. Analyses limited to complete cases implicitly assume that attrition is independent of both treatment and prognosis, an assumption rarely plausible in gerontological research. cao2022review Consequently, principled approaches that explicitly model missingness mechanisms under reasonable data assumptions are necessary to recover credible subgroup inferences. little2012prevention; seaman2013review
@@
This paper develops an integrated machine learning model tailored to pragmatic CRT that integrates mixed-effects BART into both the principal strata model and outcome models, while accounting for missing data in survival status and QoL outcomes. By augmenting BART with cluster-level random intercepts, the method preserves flexibility for nonlinear covariate effects and interactions while respecting within-cluster dependence dorie2022stan; spanbauer2021nonparametric; wundervald2022hierarchical. To accommodate incomplete follow-up common in pragmatic gerontologic trials, we adopt a nested missing at random scheme that treats missing survival status and, conditional on observed survival, missing outcomes as ignorable given treatment assignment and covariates tong2025bayesian. Within this framework we estimate SACE and CSACE and incorporate an exploratory, post-hoc “fit-the-fit” step that regresses individual-level CSACE estimates among likely always-survivors on baseline covariates to visualize effect modification. Our empirical application revisits the Whole Systems Demonstrator (WSD) telecare trial, focusing on EQ-5D–VAS at 12 months, and illustrates how conditional survivor causal effects can differ from the marginal intent-to-treat effect. The remainder of the paper presents the design and assumptions, the principal stratification estimands and missing data framework, the likelihood and mixed-effects BART specification, the WSD analysis and exploratory heterogeneity results, and concluding remarks.
@@ 2 Whole Systems Demonstrator Telecare Questionnaire Study
Whole Systems Demonstrator (WSD) Telecare Questionnaire Study henderson2013cost; hirani2014effect, a cluster-randomized trial conducted in England during 2008–2009. Randomization occurred at the general practice level, where a total of 1 189 participants across 204 practices in three local authorities (cluster sizes 1–26) were randomized to the telecare intervention or control with roughly a 1:1 ratio. The telecare intervention comprised in-home electronic sensors providing safety monitoring (e.g., fall and hazard alerts), while the control arm received usual care. Our primary analysis focuses on the QoL measured at 12 months by the EQ-5D–VAS index (range 0–100; higher values indicating a better quality of life). feng2014assessing Baseline characteristics including age, ethnicity, sex, highest level of education, number of comorbid conditions, living alone status, baseline deprivation score, physical health score, mental health score and EQ-5D–VAS index score by arm are summarized in Table 1. Baseline deprivation was measured with the English Index of Multiple Deprivation (IMD 2007), linked from participant postcodes to their Lower-layer Super Output Area; the IMD combines seven domains—income, employment, health and disability, education, crime, barriers to housing and services, and living environment—into an area-level score in which higher values indicate greater deprivation. noble2008english In WSD, IMD scores ranged from 1.89 to 66.49. Overall, characteristics were broadly comparable between arms: means of the continuous measures and the distributions of sex and ethnicity were closely aligned (all $p\geq 0.31$). The only notable difference was a modest shift in education levels (Cochran–Armitage $p=0.014$), with the intervention arm showing slightly higher proportions at the university/graduate categories.
@@
For the purposes of principal stratification, we classify a participant as a nonsurvivor when the 12 month status of is either death or being too ill to respond (due to dementia or loss of capacity, transition to long-term residential/nursing care, including sheltered housing, or data collection was fully reliant on a family caregiver). In the assembled data, 127 participants (10.7%) had outcomes truncated by death or serious deterioration; 62 (5.2%) were presumed alive but had unobserved outcomes for non–health-related reasons for example, some moved out of the area to non-participating general practice, and others no longer wished to remain in either the intervention or control group; and 237 (19.9%) had both survival status and the final outcome unobserved (e.g., prolonged hospitalization, refusal to share data, or equipment malfunction).
@@ 3.1 Notation and setup
We study a two-arm cluster-randomized trial WSD with clusters $i=1,\dots,n$, where cluster $i$ contains $N_{i}$ individuals $j=1,\dots,N_{i}$. Let $Z_{i}\in\{0,1\}$ denote the cluster-level assignment ($Z_{i}=1$ treatment; $Z_{i}=0$ control). For each individual $(i,j)$ and $z\in\{0,1\}$, $S_{ij}(z)\in\{0,1\}$ is the potential survival status at end of follow-up (11=survive, 0=death) and $Y_{ij}(z)$ is the corresponding potential non-mortality outcome, which is only well-defined when $S_{ij}(z)=1$; when $S_{ij}(z)=0$ the outcome is undefined and denoted by $\star$ zhang2003estimation. The observed data are $(Z_{i},S_{ij},Y_{ij}^{\text{obs}})$ with the usual consistency: $S_{ij}=S_{ij}(Z_{i})$ and, if $S_{ij}=1$, then $Y_{ij}^{\text{obs}}=Y_{ij}(Z_{i})$; otherwise $Y_{ij}^{\text{obs}}=\star$.
We pursue two assumptions to connect the cluster-level assignment to counterfactual outcomes. First, we invoke Stable Unit Treatment Value Assumption (SUTVA) with partial interference: for each individual $(i,j)$, the potential survival and outcome $\{S_{ij}(z),Y_{ij}(z)\}$ depend only on the assignment of their own cluster $Z_{i}$; treatment versions are unique and there is no interference across clusters, while arbitrary dependence among individuals within the same cluster is allowed. Second, we assume cluster-level randomization with positivity: for every cluster $i$, the collection $\{Y_{ij}(1),Y_{ij}(0),S_{ij}(1),S_{ij}(0)\}_{j=1}^{N_{i}}$ is independent of $Z_{i}$, and both arms occur with positive probability, $0<\Pr(Z_{i}=1)<1$. Taken together, these conditions render $Z_{i}$ ignorable for causal inference at the cluster level and underpin identification based on comparisons of observed outcomes across treatment arms.
These conditions are substantively plausible in the WSD setting. Randomization occurred at the general‐practice (cluster) level and assignment was well balanced (639 telecare vs. 550 usual care), so positivity holds and $Z_{i}$ is independent of counterfactual outcomes by design henderson2013cost; hirani2014effect. The intervention consisted of standardized, home‐installed telecare sensors delivered only to practices assigned to the telecare arm, whereas control practices continued usual care; this delivery model limits cross‐practice contamination. Any dependence among individuals within a practice (e.g., shared clinical workflows or referral patterns) is accommodated by allowing arbitrary within‐cluster correlation in our models, while interference across practices is unlikely because equipment installation and monitoring are practice‐specific and home‐based. Empirically, baseline characteristics were similar across arms (Table 1), consistent with successful randomization. Taken together, these features support SUTVA with partial interference and cluster‐level ignorability for $Z_{i}$ in the WSD trial.

### v2

@@ Abstract
Detecting heterogeneity in treatment response enriches the interpretation of gerontologic trials. In aging research, estimating the intervention’s effect on clinically meaningful outcomes poses analytical challenges when outcomes are truncated by death. For example, in the Whole Systems Demonstrator trial, a large cluster-randomized study evaluating telecare among older adults, the overall effect of the intervention on quality of life was found to be null. However, this marginal intervention estimate obscures potential heterogeneity of individuals responding to the intervention, particularly among those who survive to the end of follow-up. To explore this heterogeneity, we adopt a causal framework grounded in principal stratification, targeting the Survivor Average Causal Effect (SACE)—the treatment effect among “always-survivors,” or those who would survive regardless of treatment assignment. We extend this framework using Bayesian Additive Regression Trees (BART), a nonparametric machine learning method, to flexibly model both latent principal strata and stratum-specific potential outcomes. This enables the estimation of the Conditional SACE (CSACE), allowing us to uncover variation in treatment effects across subgroups defined by baseline characteristics. Our analysis reveals that despite the null average effect, some subgroups experience distinct quality of life benefits (or lack thereof) from telecare, highlighting opportunities for more personalized intervention strategies. This study demonstrates how embedding machine learning methods, such as BART, within a principled causal inference framework can offer deeper insights into trial data with complex features including truncation by death and clustering—key considerations in analyzing pragmatic gerontology trials.
Keywords: Whole Systems Demonstrator Telecare Questionnaire Study; gerontology trials; cluster randomized trial; pragmatic trials; principal stratification; machine learning; Bayesian inference; survivor average causal effect
@@ 1 Introduction
Pragmatic trials are a cornerstone of gerontologic research, providing rigorous yet policy-relevant evidence on interventions that aim to support independence, quality of life, and health in late life.28; 27 Their strengths lie in large, heterogeneous populations and real-world delivery, often using cluster-randomized designs. 10 However, these very characteristics can complicate the inference on the intervention performance: treatment effects averaged across diverse participants may appear modest or null, even when meaningful benefits (or harms) exist for specific subgroups. Conventional subgroup analyses, typically prespecified along a few baseline dimensions, are rarely sufficient to detect more complex forms of heterogeneity. As a result, valuable information may remain hidden, and seemingly “negative” trials risk being dismissed despite containing signals that could guide intervention targeting and future research. Therefore, machine learning and other data-driven approaches are increasingly valuable for estimating individualized treatment effects and detecting complex, multidimensional patterns of heterogeneity that conventional subgroup analyses may overlook 21; 16. As one prominent example, parallel advances in machine learning for causal inference have shown that Bayesian Additive Regression Trees (BART) can flexibly learn complex regression surfaces with built-in regularization 7, and have recently been applied to treatment effect heterogeneity detection in clinical trials. 6; 20; 13.
@@
Yet, in gerontologic trials, these opportunities come with unique complications. Quality of life (QoL) outcomes that are among the most clinically meaningful endpoints in late-life intervention studies are a prime example. 16; 25 Measures such as EQ-5D or SF-12 are widely used to capture well-being, functional status, and independence—dimensions of aging that survival alone cannot reflect. 14; 38 However, QoL outcomes face a special challenge: they are undefined for participants who die before follow-up, or who deteriorate so severely that reporting is impossible. This phenomenon, known as truncation by death, is especially common in gerontology trials, where mortality is non-negligible during extended follow-up. 30 Because participants who die may differ systematically from those who survive, and because interventions themselves may influence participant survival, analyses restricted to observed survivors risk introducing survivor bias and yield estimates that are difficult to interpret causally. One principled solution comes from principal stratification, which defines causal effects within strata of participants classified by their potential survival under both treatment and control.12; 40 In this framework, the survivor average causal effect (SACE) focuses on the “always-survivors”—those who would survive regardless of assignment—for whom both counterfactual QoL outcomes are well defined. Extending this idea to covariate conditional effects (CSACE) enables the investigation of how survivor responses vary across baseline characteristics, uncovering heterogeneity in a way that standard subgroup analyses cannot.6
@@
The death truncation challenge is compounded by the fact that many pragmatic trials in gerontology adopt cluster randomization, either at the practice, facility, or community level, to reduce contamination, facilitate implementation, or reflect the way interventions are delivered in routine care. 10; 24 Such designs are especially common when evaluating service models, long-term care interventions, or technologies that naturally operate at a group level. For example, the Whole Systems Demonstrator trial of telecare randomized at the general practice level, 33; 18 while cluster-randomized or cluster-stepped wedge designs have also been used in fall prevention, dementia care, and palliative interventions in older populations (e.g., 3; 22). Compared with individually randomized trials, CRTs introduce additional sources of correlation and heterogeneity, requiring methods that account for within-cluster dependence. A similar issue can arise in another type of closely related design in pragmatic trials called individually randomized group treatment trials (IRGTs), and a description of IRGT can be found elsewhere. 36 Recent work has estimated SACE (and covariate conditional extensions, CSACE) in CRTs using parametric principal strata and outcome models with linear mixed effects and either Bayesian or EM-based estimation. 35; 37 While these approaches are interpretable and computationally convenient, they rely on linear specifications that may miss nonlinearities and high order interactions in rich baseline covariates, and they may be sensitive to functional-form misspecification.
@@
Another common complication in pragmatic gerontologic trials is missing data. Beyond death truncation, loss to follow-up is frequent as participants relocate, withdraw consent, or become unable to complete questionnaires due to cognitive or physical decline. 23; 8 In such settings, outcomes may be missing even among survivors, creating additional uncertainty in treatment effect estimation. Moreover, survival status itself can sometimes be unobserved—for example, when participants move outside the health system or registry coverage—blurring the distinction between death-related truncation and standard missingness. 5 These challenges are especially pronounced in aging populations, where comorbidity, frailty, and mobility constraints increase the risk of incomplete follow-up 5; 1. For the estimation of treatment effect heterogeneity, failure to account for missing data can distort subgroup contrasts or mask signals of effect modification. Analyses limited to complete cases implicitly assume that attrition is independent of both treatment and prognosis, an assumption rarely plausible in gerontological research. 4 Consequently, principled approaches that explicitly model missingness mechanisms under reasonable data assumptions are necessary to recover credible subgroup inferences. 23; 31
@@
This paper develops an integrated machine learning model tailored to pragmatic CRTs that integrates mixed-effects BART into both the principal strata model and outcome models, while accounting for missing data in survival status and QoL outcomes. By augmenting BART with cluster-level random intercepts, the method preserves flexibility for nonlinear covariate effects and interactions while respecting within-cluster dependence 9; 32; 39. To accommodate incomplete follow-up common in pragmatic gerontologic trials, we adopt a nested missing at random scheme that treats missing survival status and, conditional on observed survival, missing outcomes as ignorable given treatment assignment and covariates 34. Within this framework, we estimate SACE and CSACE and incorporate an exploratory, post-hoc “fit-the-fit” step that regresses individual-level CSACE estimates among likely always-survivors on baseline covariates to visualize effect modification. Our empirical application revisits the Whole Systems Demonstrator (WSD) telecare trial, focusing on EQ-5D–VAS at 12 months, and illustrates how conditional survivor causal effects can differ from the marginal intent-to-treat effect. The remainder of the paper presents the design and assumptions, the principal stratification estimands and missing data framework, the likelihood and mixed-effects BART specification, a simulation study confirming the performance of our model and algorithm, the WSD analysis and exploratory heterogeneity results, and concluding remarks.
@@ 2 Whole Systems Demonstrator Telecare Questionnaire Study
Whole Systems Demonstrator (WSD) Telecare Questionnaire Study 17; 18 is a cluster-randomized trial conducted in England during 2008–2009. Randomization occurred at the general practice level, where a total of 1,189 participants across 204 practices in three local authorities (cluster sizes 1–26) were randomized to the telecare intervention or control with roughly a 1:1 ratio. The telecare intervention comprised in-home electronic sensors providing safety monitoring (e.g., fall and hazard alerts), while the control arm received usual care. Our primary analysis focuses on the QoL measured at 12 months by the EQ-5D–VAS index (range 0–100; higher values indicating a better quality of life). 11 Baseline characteristics including age, ethnicity, sex, highest level of education, number of comorbid conditions, living alone status, baseline deprivation score, physical health score, mental health score and EQ-5D–VAS index score by arm are summarized in Table 1. Baseline deprivation was measured with the English Index of Multiple Deprivation (IMD 2007), linked from participant postcodes to their Lower-layer Super Output Area; the IMD combines seven domains—income, employment, health and disability, education, crime, barriers to housing and services, and living environment—into an area-level score in which higher values indicate greater deprivation. 26 In WSD, IMD scores ranged from 1.89 to 66.49. Overall, characteristics were broadly comparable between arms: means of the continuous measures and the distributions of sex and ethnicity were closely aligned (all $p\geq 0.31$). The only notable difference was a modest shift in education levels (Cochran–Armitage $p=0.014$), with the intervention arm showing slightly higher proportions at the university/graduate categories.
@@
For the purposes of principal stratification, we classify a participant as a nonsurvivor when the 12 month status is either death or being too ill to respond (due to dementia or loss of capacity, transition to long-term residential/nursing care, including sheltered housing, or data collection was fully reliant on a family caregiver). In the assembled data, 127 participants (10.7%) had outcomes truncated by death or serious deterioration; 62 (5.2%) were presumed alive but had unobserved outcomes for non–health-related reasons; for example, some moved out of the area to a non-participating general practice, and others no longer wished to remain in either the intervention or control group; and 237 (19.9%) had both survival status and the final outcome unobserved (e.g., prolonged hospitalization, refusal to share data, or equipment malfunction).
@@ 3.1 Notation and setup
We study the two-arm cluster-randomized WSD trial with clusters $i=1,\dots,n$, where cluster $i$ contains $N_{i}$ individuals $j=1,\dots,N_{i}$. Let $Z_{i}\in\{0,1\}$ denote the cluster-level assignment ($Z_{i}=1$ treatment; $Z_{i}=0$ control). For each individual $(i,j)$ and $z\in\{0,1\}$, $S_{ij}(z)\in\{0,1\}$ is the potential survival status at end of follow-up (11=survive, 0=death) and $Y_{ij}(z)$ is the corresponding potential non-mortality outcome, which is only well-defined when $S_{ij}(z)=1$; when $S_{ij}(z)=0$ the outcome is undefined and denoted by $\star$ 40. The observed data are $(Z_{i},S_{ij},Y_{ij}^{\text{obs}})$ with the usual consistency: $S_{ij}=S_{ij}(Z_{i})$ and, if $S_{ij}=1$, then $Y_{ij}^{\text{obs}}=Y_{ij}(Z_{i})$; otherwise $Y_{ij}^{\text{obs}}=\star$.
We adopt two assumptions to connect the cluster-level assignment to counterfactual outcomes. First, we invoke the Stable Unit Treatment Value Assumption (SUTVA) with partial interference: for each individual $(i,j)$, the potential survival and outcome $\{S_{ij}(z),Y_{ij}(z)\}$ depend only on the assignment of their own cluster $Z_{i}$; treatment versions are unique and there is no interference across clusters, while arbitrary dependence among individuals within the same cluster is allowed. Second, we assume cluster-level randomization with positivity: for every cluster $i$, the collection $\{Y_{ij}(1),Y_{ij}(0),S_{ij}(1),S_{ij}(0)\}_{j=1}^{N_{i}}$ is independent of $Z_{i}$, and both arms occur with positive probability, $0<\Pr(Z_{i}=1)<1$. Taken together, these conditions render $Z_{i}$ ignorable for causal inference at the cluster level and underpin identification based on comparisons of observed outcomes across treatment arms.
These conditions are substantively plausible in the WSD setting. Randomization occurred at the general-practice (cluster) level and assignment was well balanced (639 telecare vs. 550 usual care), so positivity holds and $Z_{i}$ is independent of counterfactual outcomes by design. 17; 18 The intervention consisted of standardized, home-installed telecare sensors delivered only to practices assigned to the telecare arm, whereas control practices continued usual care; this delivery model limits cross-practice contamination. Any dependence among individuals within a practice (e.g., shared clinical workflows or referral patterns) is accommodated by allowing arbitrary within-cluster correlation in our models, while interference across practices is unlikely because equipment installation and monitoring are practice-specific and home-based. Empirically, baseline characteristics were similar across arms (Table 1), consistent with successful randomization. Taken together, these features support SUTVA with partial interference and cluster-level ignorability for $Z_{i}$ in the WSD trial.